At what rate must a circular space station rotate to produce a gravitation effect equal to 1/2 of the earth's gravity, g=9.8m/s^2? The ships diameter is 2.2km. Give your answer in revolutions per hour.
Answer #1

As we know

a = r w^2

9.8 / 2 = 2200* w^2

w = 0.04719* 3600 / ( 2* 3.14)

w = 27.04 rev / hr

=======

Comment before rate in case any doubt, will reply for sure.. Goodluck